# latexOverwriting file

### v1

@@ Conclusion
$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{2(q-p)}{p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
The conclusion then follows from Grönwall’s lemma. ∎
@@
The proof is similar to that of McKean’s theorem [15, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
@@
In his famous lecture notes [47], Sznitman shows for a laboratory example that
$\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [16]. Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.13),
@@ 2.5 Quantitative mean-field limit for CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s\geqslant-1$, $\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$, and consider the systems (2.8) and (2.9) with the coefficients given by
@@
Then for all $p\geqslant 2$ such that
$\frac{q}{3}\geqslant p\geqslant 2p_{\mathcal{M}}(s,\ell)=\begin{cases}2s+4,\,&\text{if $\ell=0$},\\ 2,\,&\text{if $\ell>0$},\end{cases}$
there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\leqslant CJ^{-\min\left\{\frac{p}{2},2-\frac{2p}{q}\right\}}.$
@@ 3.1 Wasserstein stability estimates for the weighted moments
In this subsection, we generalize [10, Lemma 3.2] to the more general conditions in Assumption H and, more importantly, to the weighted covariance and its square root.
@@
Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant-1$ and $u\geqslant\ell\geqslant 0$. Then:
@@
We prove first (a) and then (b). Both proofs rely on the technical result Lemma A.1 in the appendix.
@@
Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$. By [51, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$. Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [51, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$. Then
@@
Fix $R>0$. By [52, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
@@
Proceeding in the same manner, we construct a matrix $M_{\mathbf{Y}}\in\mathbf{R}^{d\times J}$ such that $\mathcal{M}_{\beta}(\nu)=M_{\mathbf{Y}}M_{\mathbf{Y}}^{\mathsf{T}}$. A result by Araki and Yamagami [3], later generalized by Kittaneh [39] and Bhatia [6], states for any two matrices $A$ and $B$ with the same shape, it holds that
@@
see also [5, Theorem VII.5.7] for a textbook presentation. This yields
@@
    When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.3 if $\ell>0$, where $f^{*}$ the supremum of $f$ over ${}^{d}$. When $\mathcal{T}=\sqrt{\mathcal{C}_{\beta}}$, the inequality (3.9) follows in the same way after noticing that
@@ 3.2 Moment estimates for the empirical measures
The result can be proved using the same reasoning as in [10, Lemma 3.4] or [23, Lemma 15]. For convenience, we denote by $C$ any constant depending only on $(f,\beta,\lambda,p,T)$. We first observe by (3.10) and Proposition A.3 that
@@
Arguing in parallel for (a) and (b), we fix $j\in$ and apply the Burkholder–Davis–Gundy inequality to find for all $t\in[0,T]$ that
@@
For the proof of the lemmas in this section, we will need [18, Theorem 1], which we recall for the reader’s convenience in Theorem 3.6 below, slightly adapted to the i.i.d. setting. We also refer to [1, Theorem 2.3], where similar results for $p=1$ and $p=2$ are proved in the context of importance sampling, with explicit constant prefactors.
@@
Since $w_{j}$ is bounded from above almost surely, it obviously holds that $\left\lVert w_{1}\right\rVert_{L^{q}(\Omega)}<\infty$. Since $\mu\in\mathcal{P}_{r}(^{d})$, it also holds that $\left\lVert V_{1}\right\rVert_{L^{r}(\Omega)}<\infty$ and $\left\lVert w_{1}V_{1}\right\rVert_{L^{s}(\Omega)}<\infty$ since $s:=\frac{pq}{q-p}<r$. Consequently, the assumptions of Theorem 3.6 are satisfied, and we obtain the statement.
@@
By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$. By an inequality due to van Hemmen and Ando [50], see also [5, Problem X.5.5], it holds that
@@ 3.4 No collapse in finite time for CBS
The evolution of the covariance matrix is governed by [11, Equation (2.18)]
@@ 4.1 Well-posedness for the microscopic models
[…]
@@
The local Lipschitz estimate on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.2) is [10, Lemma 2.1]. In order to prove the local Lipschitz continuity of $\mathcal{C}_{\beta}(\mathbf{X})$, fix $R>0$ and $J>0$. As in the proof of Proposition 3.2 we obtain
@@
Since $\mathbf{X}\mapsto\mathcal{M}_{\beta}(\mathbf{X})$ is a locally Lipschitz continuous function by [10, Lemma 2.1], it suffices to prove that $\mathbf{X}\mapsto\sqrt{w_{j}(\mathbf{X})}$ is locally Lipschitz as well. To this end, observe that
@@
By Lemma 4.1, the functions $F$ and $G$ have sublinear growth and are locally Lipschitz continuous. Thus, by the non-explosion criterion from stochastic Lyapunov theory [38, Theorem 3.5], it suffices to find a function $\phi\in C^{2}\left(\mathbf{R}^{dJ},[0,\infty)\right)$ such that
@@ 4.2 Well-posedness for the mean-field SDEs
We adapt the argument from [10], using the more general Assumption H and the improved stability estimates from Subsection 3.1. We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3. In the following, we endow the vector space $\mathcal{X}:=\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
@@
The existence of a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$ to (4.3) follows from classical theory of stochastic differential equations [45, Theorem 5.2.1]. It remains to prove the continuity statement for the weighted mean and covariance. For any $0\leqslant r\leqslant t\leqslant T$, it holds by the Burkholder–Davis–Gundy inequality that
@@
Similarly to the proof of [10, Theorem 3.1], we first construct a map
@@
where $\rho_{t}:={\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3). Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined. Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5). The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [27, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
@@
By virtue of the van Hemmen–Ando inequality [50, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
@@
which is indeed a weak formulation of the Fokker–Planck equation (2.5). For rigorous uniqueness and existence results for the Fokker–Planck equation (4.9) as well as a more precise discussion of its connection with (4.3), see for example [25]. ∎
@@
Observe that the drift coefficients of CBO and CBS are identical and that the diffusion coefficient of CBO is a globally Lipschitz continuous function of its drift coefficient. By, among other things, replacing $\sqrt{2\lambda^{-1}\mathcal{R}(\Gamma_{t})}$ in (4.3) by $S\left(\overline{X}_{t}-\mathcal{M}_{\beta}(\overline{\rho}_{t})\right)$, the above proof can be easily adapted to provide a proof for Theorem 2.3. See also [10], where a weaker version of Theorem 2.3 was already proven in the case where $p=4$ and $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$.

### v2

@@ Conclusion
$\mathbf{E}\left[\int_{0}^{t}\left\lvert b\left(X_{s}^{(j)},\mu^{J}_{s}\right)-b\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rvert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\int_{0}^{t}\left\lVert\sigma\left(X_{s}^{(j)},\mu^{J}_{s}\right)-\sigma\left(\overline{X}_{s}^{(j)},\overline{\rho}_{s}\right)\right\rVert^{p}\,\mathrm{d}s\right]\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
@@
$\mathbf{E}\left[\sup_{s\in[0,t]}\bigl{\lvert}X_{s}^{(j)}-\overline{X}_{s}^{(j)}\bigr{\rvert}^{p}\right]$ $\leqslant CJ^{-\min\left\{\frac{p}{2},\frac{q-p}{2p}\right\}}+C\int_{0}^{t}\mathbf{E}\left[\sup_{u\in[0,s]}\left\lvert X^{(j)}_{u}-\overline{X}^{(j)}_{u}\right\rvert^{p}\right]\mathrm{d}s.$
The conclusion then follows from Grönwall’s lemma, since $\frac{q-p}{2p^{2}}\leqslant\frac{q-(2\vee p_{\mathcal{M}})}{2(2\vee p_{\mathcal{M}})^{2}}$ for the value of $p$ considered. ∎
@@
The proof is similar to that of McKean’s theorem [ReviewChaintronII, Theorem 3.1], with an important difference: we eliminate a set of small probability in order to apply the local stability estimate in Item (a) on its complement.
@@
In his famous lecture notes [MR1108185], Sznitman shows for a laboratory example that
$\forall j\in\llbracket 1,J\rrbracket,\qquad\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\leqslant CJ^{-\frac{p}{2}}.$
More precisely, this statement is proved for $p=1$, and the author notes that the result generalizes to higher moment, which can be achieved by using, among other technical adaptations of his argument, the Marcinkiewicz–Zygmund inequality [MR1476912]. Although CBO does not fit within Sznitman’s simplified framework, Item (c) enables to show a bound with the same rate for the second term on the right-hand side of (2.14),
@@ 2.5 Quantitative mean-field limit for CBS
Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$ has moments of sufficiently high order. Consider the systems (2.8) and (2.9) with the coefficients given by
@@
Then for all $p>0$, there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\left\lvert X^{(j)}_{t}-\overline{X}^{(j)}_{t}\right\rvert^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\frac{1}{2}}.$ (2.19)
Remark 2.13.
In the same way as for CBO, it is possible to prove the following sharper statement. Suppose that $f\in\mathcal{A}(s,\ell,\ell)$ with $s,\ell\geqslant 0$ and that $\overline{\rho}_{0}\in\mathcal{P}_{q}(\mathbf{R}^{d})$ with $\mathcal{C}(\overline{\rho}_{0})\succ 0$. Consider the system (2.8) and (2.9) with the coefficients given by (2.18). If $q\geqslant 4p_{\mathcal{M}}(s,\ell)$, then for all $p\in(0,\frac{q}{2}]$, there is $C>0$ independent of $J$ such that
$\forall J\in\mathbf{N}^{+},\qquad\forall j\in\llbracket 1,J\rrbracket,\qquad\left(\mathbf{E}\left[\sup_{t\in[0,T]}\Big{\lvert}X^{(j)}_{t}-\overline{X}^{(j)}_{t}\Big{\rvert}^{p}\right]\right)^{\frac{1}{p}}\leqslant CJ^{-\min\left\{\frac{1}{2},\frac{q-p}{2p^{2}},\frac{q-2p_{\mathcal{M}}}{2(2p_{\mathcal{M}})^{2}}\right\}}.$ (2.20)
@@ 3.1 Wasserstein stability estimates for the weighted moments
In this subsection, we generalize [carrillo2018analytical, Lemma 3.2] to the more general conditions in Assumption H and, more importantly, to the weighted covariance and its square root.
@@
Suppose that $f\in\mathcal{A}(s,\ell,u)$ with $s\geqslant 0$ and $u\geqslant\ell\geqslant 0$. Then:
@@
The proof for both (a) and then (b) relies on the technical result Lemma A.1 in the appendix.
@@
Indeed, assume the claim holds for all such pairs of probability measures, fix $R>0$ and take any pair of probability measures $(\mu,\nu)\in\mathcal{P}_{p,R}\times\mathcal{P}_{p,R}$. By [MR2459454, Theorem 6.18], there exists a sequence $\bigl{\{}(\mu^{J},\nu^{J})\bigr{\}}_{J\in\mathbf{N}^{+}}$ in $\mathcal{P}_{p}\times\mathcal{P}_{p}$ such that $W_{p}(\mu^{J},\mu)\to 0$ and $W_{p}(\nu^{J},\nu)\to 0$ in the limit as $J\to\infty$. Since convergence in $W_{p}$ implies convergence of the moments of order up to $p$ by [MR2459454, Theorem 6.9], we can assume, discarding the beginning of the sequence if necessary, that $(\mu^{J},\nu^{J})\in\mathcal{P}_{p,2R}$ for all $J$. Then
@@
Fix $R>0$. By [MR1964483, p.5], the Wasserstein distance between empirical measures $\mu^{J}$ and $\nu^{J}$ of the form (3.6) is equal to
@@
Proceeding in the same manner, we construct a matrix $M_{\mathbf{Y}}\in\mathbf{R}^{d\times J}$ such that $\mathcal{M}_{\beta}(\nu)=M_{\mathbf{Y}}M_{\mathbf{Y}}^{\mathsf{T}}$. A result by Araki and Yamagami [Araki1981], later generalized by Kittaneh [MR0787884] and Bhatia [MR1275617], states for any two matrices $A$ and $B$ with the same shape, it holds that
@@
see also [MR1477662, Theorem VII.5.7] for a textbook presentation. This yields
@@
    When $\mathcal{T}=\mathcal{M}_{\beta}$, this follows immediately from the inequality $\operatorname{e}^{-\beta f^{*}}\leqslant\operatorname{e}^{-\beta f(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet})}\leqslant\operatorname{e}^{-\beta f_{*}}$ if $\ell=0$ and from Proposition A.3 if $\ell>0$, where $f^{*}$ the supremum of $f$ over d. When $\mathcal{T}=\sqrt{\mathcal{C}_{\beta}}$, the inequality (3.9) follows in the same way after noticing that
@@ 3.2 Moment estimates for the empirical measures
The result can be proved using the same reasoning as in [carrillo2018analytical, Lemma 3.4] or [fornasier2021consensusbased, Lemma 15]. For convenience, we denote by $C$ any constant depending only on $(f,\beta,\lambda,p,T)$. We first observe by (3.10) and Proposition A.3 that
@@
Arguing in parallel for (a) and (b), we fix $j\in\llbracket 1,J\rrbracket$ and apply the Burkholder–Davis–Gundy inequality to find for all $t\in[0,T]$ that
@@
For the proof of the lemmas in this section, we will need [MR2597592, Theorem 1], which we recall for the reader’s convenience in Theorem 3.6 below, slightly adapted to the i.i.d. setting. We also refer to [MR3696003, Theorem 2.3], where similar results for $p=1$ and $p=2$ are proved in the context of importance sampling, with explicit constant prefactors.
@@
Since $w_{j}$ is bounded from above almost surely, it obviously holds that $\left\lVert w_{1}\right\rVert_{L^{q}(\Omega)}<\infty$. Since $\mu\in\mathcal{P}_{r}(^{d})$, it also holds that $\left\lVert V_{1}\right\rVert_{L^{r}(\Omega)}<\infty$ and $\left\lVert w_{1}V_{1}\right\rVert_{L^{s}(\Omega)}<\infty$ since $s:=\frac{pq}{q-p}<r$. Consequently, the assumptions of Theorem 3.6 are satisfied, and we obtain the statement. ∎
@@
By assumption it holds that $\sqrt{\mathcal{C}_{\beta}(\overline{\mu}^{J})}\succcurlyeq\sqrt{\eta}\operatorname{I}_{d}$. By an inequality due to van Hemmen and Ando [vanHemmen1980], see also [MR1477662, Problem X.5.5], it holds that
@@ 3.4 No collapse in finite time for CBS
The evolution of the covariance matrix is governed by [CBS-Carrillo2021, Equation (2.18)]
@@ 4.1 Well-posedness for the microscopic models
$\lvert\mathcal{M}_{\beta}(\mathbf{X})\rvert\leqslant\lvert\mathbf{X}\rvert\qquad\text{and}\qquad\left\lVert\mathcal{C}_{\beta}(\mathbf{X})\right\rVert\leqslant\lvert\mathbf{X}\rvert^{2}.$ (4.1)
@@
The local Lipschitz estimate on $\mathcal{M}_{\beta}(\mathbf{X})$ in (4.2) is [carrillo2018analytical, Lemma 2.1]. In order to prove the local Lipschitz continuity of $\mathcal{C}_{\beta}(\mathbf{X})$, fix $R>0$ and $J>0$. As in the proof of Proposition 3.2 we obtain
@@
Since $\mathbf{X}\mapsto\mathcal{M}_{\beta}(\mathbf{X})$ is a locally Lipschitz continuous function by [carrillo2018analytical, Lemma 2.1], it suffices to prove that $\mathbf{X}\mapsto\sqrt{w_{j}(\mathbf{X})}$ is locally Lipschitz as well. To this end, observe that
@@
By Lemma 4.1, the functions $F$ and $G$ have sublinear growth and are locally Lipschitz continuous. Thus, by the non-explosion criterion from stochastic Lyapunov theory [khasminskii2013stochastic, Theorem 3.5], it suffices to find a function $\phi\in C^{2}\left(\mathbf{R}^{dJ},[0,\infty)\right)$ such that
@@ 4.2 Well-posedness for the mean-field SDEs
We adapt the argument from [carrillo2018analytical], using the more general Assumption H and the improved stability estimates from Subsection 3.1. We first prove Theorem 2.4, and then comment on the necessary changes for proving Theorem 2.3 in Remark 4.3. In the following, we endow the vector space $\mathcal{X}:=\mathbf{R}^{d}\times\mathbf{R}^{d\times d}$ with the norm
@@
The existence of a unique continuous strong solution $Y\in C([0,T],\mathbf{R}^{d})$ to (4.3) follows from classical theory of stochastic differential equations [MR2001996, Theorem 5.2.1]. It remains to prove the continuity statement for the weighted mean and covariance. For any $0\leqslant r\leqslant t\leqslant T$, it holds by the Burkholder–Davis–Gundy inequality that
@@
Similarly to the proof of [carrillo2018analytical, Theorem 3.1], we first construct a map
@@
where $\rho_{t}:={\rm Law}(Y_{t})$ and $(Y_{t})_{t\in[0,T]}$ is the unique solution to (4.3). Thanks to Lemma 4.2, the map $\mathcal{T}$ is well-defined. Fixed points of $\mathcal{T}$ correspond to solutions of the McKean-Vlasov SDE (1.5). The existence of fixed points follows from applying the Leray-Schauder Fixed Point Theorem [gilbarg1977elliptic, Chapter 11] in the space $C\left([0,T],\mathcal{X}\right)$, once we have proved that $\mathcal{T}$ is compact and that the following set is bounded:
@@
By virtue of the van Hemmen–Ando inequality [vanHemmen1980, Proposition 3.2], it holds for any unitarily invariant norm, thus in particular for the Frobenius norm, that
@@
which is indeed a weak formulation of the Fokker–Planck equation (2.5). For rigorous uniqueness and existence results for the Fokker–Planck equation (4.9) as well as a more precise discussion of its connection with (4.3), see for example [MR2830391]. ∎
@@
Observe that the drift coefficients of CBO and CBS are identical and that the diffusion coefficient of CBO is a globally Lipschitz continuous function of its drift coefficient. By, among other things, replacing $\sqrt{2\lambda^{-1}\mathcal{R}(\Gamma_{t})}$ in (4.3) by $S\left(\overline{X}_{t}-\mathcal{M}_{\beta}(\overline{\rho}_{t})\right)$, the above proof can be easily adapted to provide a proof for Theorem 2.3. See also [carrillo2018analytical], where a weaker version of Theorem 2.3 was already proven in the case where $p=4$ and $f\in\mathcal{A}(1,0,0)\cup\mathcal{A}(1,2,2)$.